In a historical movie, two knights on horseback start from rest 64 m apart and ride directly toward each other to do battle. Sir George's acceleration has a magnitude of 0.21 m/s2, while Sir Alfred's has a magnitude of 0.29 m/s2. Relative to Sir George's starting point, where do the knights collide?
Answer #1

DATA

:

б4 772

aG0.21 m/s2

A = 0.29 m/s2

x=?

:

SOLUTION

:

Set origin at Sir George's starting point. The position of Sir George's as a function of time can be expressed as

:

cL (1)

:

The position of Sir Alfred's as a function of time can be expressed as

:

aat (2) LA

:

Let tc be the collision time, then

:

x_{G}(t=t_{c})=x_{A}(t=t_{c})equiv x

:

From eq. (1),

:

x_{G}(t=t_{c})=x=rac{a_{G}t_{c}^{2}}{2}

te 2

:

From eq. (2),

:

x_{A}(t=t_{c})=x=x_{oA}-rac{a_{A}t_{c}^{2}}{2} : : : : : : : : : : (4)

:

inserting eq. (3) into (4), we obtain

:

x=x_{oA}-rac{a_{A}}{2}left ( rac{2x}{a_{G}} ight )

Rightarrow : : : x=x_{oA}- rac{xa_{A}}{a_{G}}

TCA aG

Rightarrow : : : xleft (rac{a_{G}+a_{A}}{a_{G}} ight )=x_{oA}

Rightarrow : : : x=rac{x_{oA}a_{G}}{a_{G}+a_{A}}

Rightarrow : : : x=rac{(64m)(0.21 , m/s^{2})}{0.21 , m/s^{2}+0.29, m/s^{2}}

:

or

:

{color{Blue} x=26.9 , m}

:

The knights collide 26.9 m away from the Sir George's starting point.
Answer #2

SOLUTION :


Let they collide after t secs. 


So,


Distance travelled by Sir George = 1/2 a t^2 = 1/2 * 0.21 * t^2 = 0.105 t^2


Distance travelled by Sir Alfred = 1/2 a’ t^2 = 1/2 * 0.29* t^2 = 0.145 t^2


As they collide, the sum of their distances travelled = 91 m

=> 0.105 t^2 + 0.145 t^2 = 64

=> 0.250 t^2 = 64

=> t^2 = 64 / 0.25 = 256 


Sir George ’s distance travelled = 0.105 * 256 = 26.88 m 


Hence,  they collide at a  point 26.88  m  from the starting point of Sir George. (ANSWER).

answered by: Tulsiram Garg

> In the fifth line, please correct , " 91 m " as " 64 m "

Tulsiram Garg Sun, Oct 10, 2021 3:17 PM